OpenStudy (anonymous):

Find the volume of a cylinder. 12 cm- diameter 50 cm (height) Round your answer to the nearest cubic centimeter. Please help me with this question.

OpenStudy (anonymous):

yeah. I guess

OpenStudy (anonymous):

radius = diameter/2 r = 12/2 = 6 V=πr²h V=π(6)²(50) V=5655cm^3 (rounded)